Circular No. 6282

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

X-RAY NOVA 1992 IN SMALL MAGELLANIC CLOUD G. Clark and R. Remillard, Massachusetts Institute of Technology; and

J. Woo, Harvard-Smithsonian Center for Astrophysics, report: “A ROSAT observation of 1992 Oct. 1–2 discovered an x-ray nova located 5′ southeast of SMC X-1 at α = 1^{h}17^{m}41^{s}.9, δ = −73^{o}31′01′′ (equinox 2000.0). The object was not detected in a similar observation 359 days earlier, but it was detected 246 days later at a counting rate diminished by a factor of 267, which corresponds to an average e-folding decay time of 44 days. The flux varied on a timescale of an hour or less during the 1992 observation, but no periodic pulsations have been detected. The 1992 flux level would be produced by a source in the SMC radiating isotropically with a luminosity in excess of 2.5 × 10^{37} erg/s. The 5′′-radius positional error circle includes a blue star of mag ∼ 14 that should be an easy target for the optical measurements required to determine whether it is the binary companion of a blackhole. If it is, this nova will be the first blackhole x-ray nova with a massive early-type companion.”

COMET C/1995 Q1 (BRADFIELD) Continuation to the ephemeris on IAUC 6225 (elements MPC 26143):

1995 /96 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 29 2^{h}09.^{m}22 +65^{o}35.^{′}7 1.640 2.309 121.^{o}3 21.^{o}3 10.7
Jan. 8 1 55.61 +59 15.2 1.876 2.451 114.4 21.4 11.3
18 1 52.96 +54 24.1 2.138 2.590 106.1 21.4 11.8
28 1 55.35 +50 47.0 2.416 2.726  97.5 21.0 12.3
Feb. 7 2 00.43 +48 07.3 2.702 2.859  88.9 20.2 12.7
17 2 07.10 +46 11.3 2.989 2.990  80.5 19.0 13.1
27 2 14.77 +44 48.7 3.272 3.118  72.3 17.6 13.5
Mar. 8 2 23.06 +43 51.6 3.547 3.244  64.4 16.0 13.9
18 2 31.76 +43 14.3 3.808 3.369  56.8 14.3 14.2
28 2 40.69 +42 52.5 4.055 3.491  49.5 12.6 14.5
Apr. 7 2 49.70 +42 42.9 4.283 3.611  42.7 10.8 14.7
17 2 58.71 +42 43.1 4.491 3.730  36.3  9.2 15.0
27 3 07.60 +42 51.3 4.676 3.847  30.8  7.7 15.2

Visual magnitudes: Nov. 16.12 UT, 9.6 (C. E. Spratt, Victoria, BC, 0.20-m refl.); 18.03, 9.4 (J. D. Shanklin, Cambridge, England, 14 × 100 binoculars); 18.31, 10.6 (A. Hale, Cloudcroft, NM, 0.41-m refl.); 20.12, 9.9 (Spratt); 30.49, 11.0 (Hale); Dec. 23.12, 10.1 (Spratt); 24.12, 10.5 (Spratt).

1996 January 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden